OpenStudy (s):

How to find two positive x and y that minimize Q=x^2y if x+y+2 ?

OpenStudy (anonymous):

did u mean x+y=2?

OpenStudy (s):

oh, yes, sorry x+y=2

OpenStudy (anonymous):

x+y = 2 y = 2-x Q=(x^2)*y = x^2(2-x) = 2x^2-x^3 Q' = 4x - 3x^2 = 0 x(4-3x)=0 x=0; 4-3x=0, -3x=-4; x=4/3

OpenStudy (anonymous):

@S did you leave? Because it looks like @Brutal was putting in some time to help you here...

OpenStudy (s):

no , i am steel here, just trying to understand this

OpenStudy (s):

still*

OpenStudy (anonymous):

Cool, no worries then - as you were ;)

OpenStudy (anonymous):

okay just keep the positive positive value: x=4/3 x=4/3, but you know that y = 2- x, so y = 2-(4/3) = (6-4)/3 = 2/3

OpenStudy (s):

yes, was just thinking about y value =D, thank you soo much for your help!

OpenStudy (anonymous):

Great job both of you @S and @Brutal. Thank you from OpenStudy!

OpenStudy (anonymous):

np :)